What is the probability of drawing a red 9?
Since there are 2 red nines in a 52 card deck, the probability would be 2/52 or 1/26.
How do you find the probability of picking a king of spades and the ace of hearts out of 52 cards?
2 in 2652, or about 0.0007541. (2 in 52 times 1 in 51)
Detailed Explanation:
Initially, you have 52 cards. The King of Spades and the Ace of Hearts are two of them, so you have a 2 in 52 chance of picking the King of Spades or the Ace of Hearts on your first draw. On your second draw, you have to pick the other card, but now there are only 51, so the chance of picking the other one is 1 in 51.
Since these are independent, though sequentially related events, to compute the probability that both events occur, you simply multiply the two probabilities together. 2 in 52 times 1 in 51 is 2 in 2652, or about 0.0007541.
2
What is the probability of having twin calves?
about 1-4%.
Animal twins
Twins are common in many animal species, such as cats, sheep, ferrets and deer. The incidence of twinning among cattle is about 1-4%, and research is under way to improve the odds of twinning, which can be more profitable for the breeder if complications can be sidestepped or managed. The nine-banded armadillo (Dasypus novemcinctus) has identical twins (usually four babies) as its regular reproduction and not as exceptional cases.
Twin calves of the Hereford breed
I hope this can help you.
What is the probability of getting tails if a coin is tossed 4 times?
If you look at the as the probability of getting 1 or more tail in 4 coin tosses, you would then calculate the probability of tossing 4 heads in a row and subracting that from 1. The probability fo tossing 4 heads is 1/2 * 1/2 * 1/2 * 1/2 = 1/16. 1 - 1/16 = 15/16.
It is 0.2549, approx.
What is the probability of getting two heads?
If you are talking about coins, 25%
If you are talking about genetics, then it is less than 1%.
What is the probability tossing 3 nickels?
Depends on how many nickels you have. If you have less than three nickels the probability is zero. If not, 100% (should you decide to throw them).
How many 10's are in a deck of 52 cards?
4
a 10 of hearts, a 10 of spades, a 10 of clubs, and a 10 of diamonds
What is the probability of rolling a number that is a multiple of 3 on a fair number cube?
A fair cube has six faces numbered 1, 2, 3, 4, 5, 6
3 and 6 are multiples of 3
so you have 2 chances in 6
or 1 in 3 in it's simplest form
What considerations determine physical distribution of a product?
manufacturers need to review issues such as distribution objectives, product transportation, and product warehousing. Choosing the mode of transportation requires an understanding of each possible method:
What is the probability of drawing a black card in a standard deck?
Excluding jokers, the probability is 1 in 2.
What is the probability of getting a 7 after rolling a single die numbered 1to 6?
What do you think? Its a tiny bit very obvious.
What is the probability for family to born 2 boys and a girl?
There is no simple answer to the question because the children's genders are not independent events. They depend on the parents' ages and their genes.
However, if you assume that they are independent events then, given that the probability of a boy is approx 0.52, the probability of the other two being boys is 0.4994
What is the probability of rolling first an odd number then an even number with one fair die?
Prob = 1/2 * 1/2 = 1/4
What is the relevance of calculating standard deviation?
The Standard Deviation will give you an idea of how 'spread apart' the data is. Suppose the average gasoline prices in your town are 2.75 per gallon. A low standard deviation means many of the gas stations will have prices close to that price, while a high standard deviation means you would find prices much higher and also much lower than that average price.
The probability that a single person would like at least ONE flavour - is 9/10
* * * * *
No.
350 liked only Vanilla
250 liked Vanialla and Chocolate
50 liked only Chocolate
That makes 650 [the remaining 350 did not like either].
Therefore the probability that a randomly selected person likes at least one of the two tastes is 650/1000 or 65%
What is the probability of choosing an even number out of seven counters?
So I guess the counters are numbered 1 through 7, and you pick one of them.
There are 7 possibilities, 3 of them are even numbers {2,4,6}, so the probability is 3 out of 7. 3/7 = approx. 42.86%
Square centimeter, a unit of area: the area equivalent to a square that has a side length of 1 cm.